OpenStudy (anonymous):

In y=mx + b form state the equation that goes through the points (0, 6) and (2, 14). (Hint: Find the slope and state the value of the y-intercept.) Medals will be rewarded! :p

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (the_fizicx99):

Ok to find the slope you use the formula: \[M = \frac{ Y2 - Y1 }{ X2-X1 }\]

OpenStudy (the_fizicx99):

So plug in

OpenStudy (anonymous):

so m = 6 - 14/ 0 - 2

OpenStudy (anonymous):

m = 6^2 - 14^2/0^2 - 2^2

OpenStudy (the_fizicx99):

No no, Y2 - y1, your second y. \[m = \frac{ 14 - 6 }{ ? }\] you do the x's

OpenStudy (anonymous):

2 - 0

OpenStudy (anonymous):

m = 14 - 6/2 - 0

OpenStudy (the_fizicx99):

Yes so (0, 6) and (2, 14) \[M = \frac{ 14 - 6 }{ 2 - 0 }\] Simplify now

OpenStudy (anonymous):

ok m = 8/2

OpenStudy (the_fizicx99):

Ok, that can be further simplified. 8 / 2 = ?

OpenStudy (anonymous):

4

OpenStudy (the_fizicx99):

Yes! So M = 4 Your slope or M = 4 Now you plug in: Y - y1 = M(x-x1)

OpenStudy (anonymous):

?, ok now I'm confused do I have to make an equation

OpenStudy (the_fizicx99):

No no, you plug in your values into that formula, what is your y1 ?

OpenStudy (anonymous):

14

OpenStudy (the_fizicx99):

Nope, look again, that is your y2. Y1 the first ordered pair. (x,y)

OpenStudy (anonymous):

6?

OpenStudy (the_fizicx99):

Yup!

OpenStudy (the_fizicx99):

Ok, so what's your M ?

OpenStudy (the_fizicx99):

M = slope btw

OpenStudy (anonymous):

4

OpenStudy (the_fizicx99):

And x1 ?

OpenStudy (anonymous):

0

OpenStudy (the_fizicx99):

kk, Y1 = 6 M = 4 x1 = 0 Now plug those values into: Y-y1 = M(x-x1)

OpenStudy (anonymous):

Y - 6 = 4(x - 0)

OpenStudy (anonymous):

Is this the asnwer

OpenStudy (the_fizicx99):

Yup, now simplify. First multiply 4*x and 4*0

OpenStudy (the_fizicx99):

No there's a bit more to it :P

OpenStudy (anonymous):

4x and 0

OpenStudy (the_fizicx99):

Now rewrite it, y-6 =4x + 0 Add 6 to both sides to have y = mx + b

OpenStudy (the_fizicx99):

Now keep in mind that in Y = Mx + B The M is the slope, and the B is the y intercept.

OpenStudy (anonymous):

Y= 6? Mx= 4 B= ?

OpenStudy (the_fizicx99):

No no, first rewrite it. y-6=4x + 0 Add 6 to both sides. y-6 = 4x + 0 +6 +6 y = 4x + 6

OpenStudy (the_fizicx99):

Now you have it in form of y = mx + b Y = 4x + 6 4 = M The y intercept is B B = 6

OpenStudy (anonymous):

Oh, now I get this makes a lot more sense thx for your help. I actually have other question if you would like to help me on them

OpenStudy (the_fizicx99):

Np!